# Almost sure convergence rates of stochastic gradient methods under gradient domination

### v1

@@ 1 Introduction
(i) Under global gradient domination with parameter $\beta$ (on the entire function domain), we prove that the last iterate of stochastic gradient descent (SGD) and stochastic heavy ball (SHB) converge almost surely and in expectation towards the global optimum with rate arbitrarily close to $o(n^{-\frac{1}{4\beta-1}})$. The almost sure and expectation rates of convergence that we obtain depend on the gradient domination parameter $\beta$ and are the same for both algorithms and convergence types. For SGD this rate is arbitrarily close to the tight upper bound known in expectation [12], while the almost sure convergence rate is new for the (weak) gradient domination assumption (see Theorem 4.1 and discussion afterwards). To the best of our knowledge for SHB this is the first convergence result towards global optima under (weak) gradient domination, for both almost sure convergence and convergence in expectation (see Theorem 4.2).
(ii) We consider the case where the gradient domination property holds only locally, either around stationary points or around global minima. We provide the first local convergence rates under these settings: we prove that SGD remains within the good local region with high probability and, conditioned on this event, we obtain converges rates almost surely and in expectation (see Theorem 5.1 and Theorem 5.2).
(iii) Our local setting covers generic classes of functions. In particular, we demonstrate that it encompasses the training task of deep neural networks with analytic activation functions in supervised learning. Our result illustrates that the iterates of SGD are likely to become trapped in areas of local minima when the step size is is small. We verify under mild conditions, that SGD converges to local minima with given convergence speed (see Corollary 6.1).
(iv) Finally, we derive novel convergence results for policy gradient methods in reinforcement learning. We show that local gradient domination holds around the global optimum for the softmax parametrization resulting in the first local convergence rate for stochastic policy gradient with arbitrary batch-size (see Corollary 7.2).
We summarize the contributions of this paper in Table 1. These findings are also illustrated in a numerical toy experiment in Appendix E, where we have implemented SGD and SHB for monomials with increasing degree.
@@
$\frac{1}{2}$ $\Theta\left(n^{-1+\epsilon}\right)$ $o\left(n^{-1+\epsilon}\right)$ global SGD a.s. 4.1 (i); e.g. [27, Thm. 1]
$\mathbb{E}$ 4.1 (ii); e.g. [20, Thm. 3]
SHB a.s 4.2 (i); e.g. [27, Thm. 2]
$\mathbb{E}$ 4.2 (ii); e.g. [25, Thm. 4.3]
local* SGD a.s. 5.1 (ii); 5.2 (ii)
$\mathbb{E}$ 5.1 (iii); 5.2 (iii); e.g. [35, Thm. 4]
@@
$(\frac{1}{2},1]$ $\Theta\left(n^{-\frac{2\beta}{4\beta-1}}\right)$ $o\left(n^{-\frac{1}{4\beta-1}+\epsilon}\right)$ global SGD a.s. 4.1 (i)
$\mathbb{E}$ 4.1 (ii); e.g. [12, Cor. 1]
SHB a.s. 4.2 (i)
$\mathbb{E}$ 4.2 (ii)
local* SGD a.s. 5.1 (ii); 5.2 (ii)
$\mathbb{E}$ 5.1 (iii); 5.2 (iii)
Related work.
We give a brief overview of related literature. For a more detailed literature review and classification of our contributions we refer the reader to Section A.1.
@@
Usually, convergence rates for stochastic gradient methods are analysed in expectation or with high-probability bounds. Though originally, motivated by [46], research commenced with the quest for almost sure convergence rates for gradient methods. In recent years, almost sure convergence rates towards global optima have been derived under (strong) convexity [49, 27]. The authors in [49] also analyzed almost sure convergence rates for SHB but under the assumption of convexity and [27] studied SHB under (strong) convexity and in a non-convex setting. Under the global gradient domination property almost sure convergence to a critical point is provided in [7], though without a rate.
There has been a lot of effort to derive local convergence guarantees for stochastic first order methods. It was shown in [35] that, subject to certain assumptions on the objective function, the SGD scheme converges almost surely towards a local minimum. By assuming strong convexity within a neighbourhood of a local minimum, they further established a local convergence rate conditioned on the event of remaining in this neighbourhood. We extend ideas from the strongly convex setting to local gradient domination and derive local convergence rates, both in the almost sure and expectation sense. In our analysis we distinguish the cases where the local gradient domination property holds in a neighbourhood of a local minimum or in the neighbourhood of the global optimum respectively.
@@
For the application in the training of DNNs, it is worth noting that local convergence of SGD has been analyzed under stronger variants of gradient domination [52, 1]. Due to the stronger form of gradient domination, specific sub-classes of DNNs need to be considered to verify these assumptions whereas our result is only constrained to analytic activation functions.
@@
For the application in reinfordement learning, recent results showed that choosing the tabular softmax parametrization in policy gradient (PG) algorithms results in objective functions which fulfill a non-uniform gradient domination property [34, 55, 21]. While convergence of PG for exact gradients is well understood, convergence rates for stochastic PG are rare and mostly require very large batch sizes [9, 21, 10]. It is noteworthy that a similar local analysis for stochastic policy gradient under entropy regularization is presented in [10]. Their local result is also based on [35], but requires an increasing batch size sequence to obtain $O(\frac{1}{n})$-convergence towards the regularized optimum with high probability. In contrast, we consider both the unregularized and entropy regularized setting and observe that one can also achieve convergence arbitrarily close to $o(\frac{1}{n})$ without the need for an increasing batch size. Moreover, the local convergence occurs almost surely on an event with high probability.
@@ 2 Mathematical Background - Optimization under Gradient Domination
Applying (2) and the iteration scheme yields the iterative descent property
@@
with exponent $\beta=1/2$ [45]. It is worthwhile to emphasize that (3) is weaker than strong convexity, the classical textbook assumption that fails for many applications. Under the PL-condition the iterative descent property can be written as a recursion:
@@
    Let $x^{\ast}\in\mathbb{R}^{d}$ be a stationary point, i.e. $\nabla f(x^{\ast})=0$. We say that $f$ satisfies a local gradient domination property in $x^{\ast}$ with parameter $\beta_{x^{\ast}}\in[\frac{1}{2},1]$ if there exist a radius $r_{x^{\ast}}>0$ and a constant $c_{x^{\ast}}>0$ such that
    $\|\nabla f(x)\|\geq c_{x^{\ast}}|f(x)-f(x^{\ast})|^{\beta_{x^{\ast}}}$
@@
The PL-condition mentioned above is a special case of the general global gradient domination property for $\beta=\frac{1}{2}$. In [30] it has been demonstrated that all analytic functions satisfy the local gradient domination property, emphasizing the particular significance of the local case. Further, it has been proved that all overparametrized neural networks fulfill the local gradient domination property [26]. See also [31, 8, 15] and references therein for the application of (strong) gradient domination to (deep) neural networks. In [12, 2, 4, 59] examples of functions are discussed that fulfill the (weak) gradient domination property. For instance, one-dimensional monomials $f(x)=|x|^{p}$, $p\geq 2$, satisfy the weak global gradient domination property with $\beta=\frac{p-1}{p}$. We refer to [12, Appendix A] for a longer list of globally gradient dominated functions including convex and non-convex functions. Notably, in reinforcement learning it is known that the tabular softmax parametrization leads to a parametrized value function that satisfies the so-called ”non-uniform” PL-inequality [34, 33, 21]. In Section 7 we will show how this non-uniform gradient domination implies local gradient domination for $f^{*}$. This renders our local analysis of stochastic gradient methods specifically applicable in RL. As mentioned earlier, since every analytic function already satisfies local gradient domination, we expect that the local analysis can encompass further parametrizations, such as neural networks.
Assumptions on the Stochastic First Order Oracle.
@@
It is worth noting that the (ABC) assumption is a generalization of the bounded variance assumption that appears for $A=B=0$. It was introduced by [20] as expected smoothness condition and shown to be the weakest assumption among many others.
Stochastic Gradient Methods.
The following two classical optimization algorithms will be analyzed in this article. Both algorithms are described as discrete time stochastic process $(X_{n})_{n\in\mathbb{N}}$ driven by noisy gradient evaluations in (4). In each iteration, we assume that the stochastic first order oracle is accessed through the evaluation of $\zeta_{n+1}$ which is a copy of $\zeta$ independent from the current state $X_{n}$.

### v3

@@ 1 Introduction
  1. (i)
    Under global gradient domination with parameter $\beta$ (on the entire function domain), we prove that the last iterate of stochastic gradient descent (SGD) and stochastic heavy ball (SHB) converge almost surely and in expectation towards the global optimum with rate arbitrarily close to $o(n^{-\frac{1}{4\beta-1}})$. The almost sure and expectation rates of convergence that we obtain depend on the gradient domination parameter $\beta$ and are the same for both algorithms and convergence types. For SGD this rate is arbitrarily close to the tight upper bound known in expectation [11], while the almost sure convergence rate is new for the (weak) gradient domination assumption (see Theorem 4.1 and discussion afterwards). To the best of our knowledge for SHB this is the first convergence result towards global optima under (weak) gradient domination, for both almost sure convergence and convergence in expectation (see Theorem 4.2).
  2. (ii)
    We consider the case where the gradient domination property holds only locally, either around stationary points or around global minima. We provide the first local convergence rates under these settings: we prove that SGD remains within the good local region with high probability and, conditioned on this event, we obtain converges rates almost surely and in expectation (see Theorem 5.1 and Theorem 5.2).
  3. (iii)
    Our local setting covers generic classes of functions. In particular, we demonstrate that it encompasses the training task of deep neural networks with analytic activation functions in supervised learning. Our result illustrates that the iterates of SGD are likely to become trapped in areas of local minima when the step size is small. We verify under mild conditions, that SGD converges to local minima with given convergence speed (see Corollary 6.1).
  4. (iv)
    Finally, we derive novel convergence results for policy gradient methods in reinforcement learning. We show that local gradient domination holds around the global optimum for the softmax parametrization resulting in the first local convergence rate for stochastic policy gradient with arbitrary batch-size (see Corollary 7.2).
We summarize the contributions of this paper in Table 1. These findings are also illustrated in a numerical toy experiment in Appendix B, where we have implemented SGD and SHB for monomials with increasing degree.
@@
$\frac{1}{2}$ $\Theta\left(n^{-1+\epsilon}\right)$ $o\left(n^{-1+\epsilon}\right)$ global SGD a.s. Theorem 4.1 (i);
Liu and Yuan [28, Thm. 1]
$\mathbb{E}$ Theorem 4.1 (ii);
Khaled and Richtárik [21, Thm. 3]
SHB a.s. Theorem 4.2 (i);
Liu and Yuan [28, Thm. 2]
$\mathbb{E}$ Theorem 4.2 (ii);
Liang et al. [26, Thm. 4.3]
local* SGD a.s. Theorem 5.1 (ii);
Theorem 5.2 (ii)
$\mathbb{E}$ Theorem 5.1 (iii); Theorem 5.2 (iii);
Mertikopoulos et al. [38, Thm. 4]
@@
$(\frac{1}{2},1]$ $\Theta\left(n^{-\frac{2\beta}{4\beta-1}}\right)$ $o\left(n^{-\frac{1}{4\beta-1}+\epsilon}\right)$ global SGD a.s. Theorem 4.1 (i)
$\mathbb{E}$ Theorem 4.1 (ii);
Fatkhullin et al. [11, Cor. 1]
SHB a.s. Theorem 4.2 (i)
$\mathbb{E}$ Theorem 4.2 (ii)
local* SGD a.s. Theorem 5.1 (ii);
Theorem 5.2 (ii)
$\mathbb{E}$ Theorem 5.1 (iii);
Theorem 5.2 (iii)
1.1 Literature Review and Classification of our Contribution
The roots of stochastic gradient methods trace back to Robbins and Monro [51]. Since then, various variants of SGD have been established as fundamental algorithms for optimizing complex models in the realm of machine learning. We refer to Bottou et al. [5] and Garrigos and Gower [17] for a detailed overview.
We start the review with the literature deriving convergence rates in expectation for SGD. Under the assumptions of smoothness and (strong) convexity Polyak [47], Moulines and Bach [39], Nguyen et al. [43], Wang et al. [54], Liu et al. [30] studied convergence rates towards global optima. Moreover, many articles additionally analyze the non-convex case and prove convergence rates for the gradient norm towards zero [19, 25, 30, 44].
Notably, several other results regarding convergence of SGD towards global optima have been established under the gradient domination setting [20]. Bassily et al. [3], Liu et al. [27] demonstrate exponential convergence rates in expectation in the overparameterized setting under strong gradient domination. See also Madden et al. [33], where high-probability convergence bounds are shown using the (strong) gradient domination property. Further high-probability bounds on the approximation error are provided in Scaman et al. [52] under a generalized gradient domination property, the so-called Separable-Łojasiewicz assumption, fulfilled by smooth neural networks. In Lei et al. [24] also strong gradient domination is assumed, where the smoothness assumption is weakened through $\alpha$-Hölder continuity, achieving a rate of $O(\frac{1}{n^{\alpha}})$ in expectation. The Equation ABC condition is introduced in Khaled and Richtárik [21] and where $O(\frac{1}{n})$ convergence is shown under strong gradient domination. Furthermore, Fatkhullin et al. [11] and Fontaine et al. [14] consider generalizations of gradient domination that include our definition as a special case. They derive convergence rates in expectation which we encompass with our result and extend to almost sure convergence (see also the discussion behind Theorem 4.1). Finally, Masiha et al. [35] established both upper and lower bounds for projected gradient descent under gradient domination.
@@ 1.1 Literature Review and Classification of our Contribution
All the results mentioned so far consider convergence in expectation or high-probability bounds, although originally, motivated by Robbins and Siegmund [50], research commenced with the quest for almost sure convergence rates for gradient methods. In recent years, Sebbouh et al. [53] and, building upon it, Liu and Yuan [28] derive almost sure convergence rates towards global optima under convexity for SGD and SHB. In addition, Liu and Yuan [28] study SHB and Nesterov acceleration under PL. Returning the attention back to SGD with respect to gradient domination also some almost sure convergence results have been established. As an extension to the PL-type gradient domination, in Chouzenoux et al. [7] the so-called KL property is assumed, which contains gradient domination as a special case. The authors demonstrate almost sure convergence to a critical point, though without a rate. To conclude, to the best of our knowledge the derived almost sure convergence rate under gradient domination in Theorem 4.1 is novel.
Next, we want to provide further insights to the literature regarding SHB. In the realm of momentum methods, Polyak’s Heavy-Ball Method (HBM) [46] and Nesterov’s accelerated gradient method [42] stand out as a foundational contribution. The authors of Gadat et al. [16] provide a detailed description of the stochastic formulation of HBM and establish almost sure convergence but without giving a rate. In Yang et al. [58], Orvieto et al. [45], Yan et al. [57], Mai and Johansson [34], Zhou et al. [62] convergence rates in expectation are shown in (strongly) convex and non-convex settings, where the non-convex analysis covers convergence of the norm of the gradient. Convergence of momentum methods under the strong gradient domination property and linear convergence due to an overparametrized machine learning setting is shown in Gess and Kassing [18]. Liang et al. [26] determine $O(\frac{1}{n})$ convergence rate for SHB under strong gradient domination. Our main result for SHB presented in Theorem 4.2 describes almost sure convergence and convergence in expectation under global gradient domination. Both result are quantified with a given rate of convergence.
In the following, we aim to differentiate our local convergence analysis from existing results in the literature. There has been a lot of effort to derive local convergence guarantees for stochastic first order methods. In Dereich and Kassing [8] almost sure convergence of SGD to a stationary point under the local gradient domination (for $x^{\ast}$) is demonstrated, provided that the process ($X_{n}$) remains local, albeit without a rate. A local analysis of SGD towards minima without any gradient domination assumption is presented in Fehrman et al. [13]. Instead, a rank assumption is imposed on the Hessian, and mini-batches, along with resampling, are leveraged to ensure convergence to the global optimum with high probability. The resulting rate does not converge to zero and requires an increasing batch size. Finally, under the global Lipschitz assumption on the objective, it can be verified that SGD almost surely converges to a stationary point as demonstrated by Mertikopoulos et al. [38]. In the same work the authors derive a local convergence analysis under local strong convexity. Our analysis in Section 5 builds upon Mertikopoulos et al. [38] and generalizes their results to the local gradient domination property. In our analysis we distinguish the cases where the local gradient domination property holds in a neighbourhood of a local minimum or in the neighbourhood of the global optimum respectively. Finally, we would like to acknowledge that related results have been independently obtained in the recent preprint [49].
@@
For the application in the training of DNNs, it is worth noting that local convergence of SGD has been analyzed under stronger variants of gradient domination by [56, 1]. Due to the stronger form of gradient domination, specific sub-classes of DNNs need to be considered to verify these assumptions whereas our result is only constrained to analytic activation functions. Under the machine learning noise conditions in [56], convergence toward zero loss with high probability is shown, provided that the initial loss is sufficiently small. In contrast, [1] demonstrate convergence towards zero loss under initialization in a local (strong) Łojasiewicz region. Indeed, one can construct DNNs satisfying the latter condition [6].
@@
For the application in reinforcement learning, recent results showed that choosing the tabular softmax parametrization in policy gradient (PG) algorithms results in objective functions which fulfill a non-uniform gradient domination property [36, 59, 22]. While convergence of PG for exact gradients is well understood, convergence rates for stochastic PG are rare and mostly require very large batch sizes [9, 22, 10]. It is noteworthy that a similar local analysis for stochastic policy gradient under entropy regularization is presented in Ding et al. [10]. Their local result is also based on Mertikopoulos et al. [38], but requires an increasing batch size sequence to obtain $O(\frac{1}{n})$-convergence towards the regularized optimum with high probability. In contrast, we consider both the unregularized and entropy regularized setting and observe that one can also achieve convergence arbitrarily close to $o(\frac{1}{n})$ without the need for an increasing batch size. Moreover, the local convergence occurs almost surely on an event with high probability.
@@ 2 Mathematical Background - Optimization under Gradient Domination
Applying Equation 2 and the iteration scheme yields the iterative descent property
@@
with exponent $\beta=1/2$ [48]. It is worthwhile to emphasize that Equation 3 is weaker than strong convexity, the classical textbook assumption that fails for many applications. Under the PL-condition the iterative descent property can be written as a recursion:
@@
    Let $x^{\ast}\in\mathbb{R}^{d}$ be a stationary point, i.e. $\nabla f(x^{\ast})=0$. We say that $f$ satisfies a local gradient domination property in $x^{\ast}$ with parameter $\beta_{x^{\ast}}\in[\frac{1}{2},1]$ if a radius $r_{x^{\ast}}>0$ and a constant $c_{x^{\ast}}>0$ such that
    $\|\nabla f(x)\|\geq c_{x^{\ast}}|f(x)-f(x^{\ast})|^{\beta_{x^{\ast}}}$ (4)
@@
The PL-condition mentioned above is a special case of the general global gradient domination property for $\beta=\frac{1}{2}$. In Lojasiewicz [32] it has been demonstrated that all analytic functions satisfy the local gradient domination property, emphasizing the particular significance of the local case. Further, it has been proved that all overparametrized neural networks fulfill the local gradient domination property [27]. See also Madden et al. [33], Dereich and Kassing [8], Frei and Gu [15] and references therein for the application of (strong) gradient domination to (deep) neural networks. In Fatkhullin et al. [11], Attouch et al. [2], Bolte et al. [4], Zhou et al. [63] examples of functions are discussed that fulfill the (weak) gradient domination property. For instance, one-dimensional monomials $f(x)=|x|^{p}$, $p\geq 2$, satisfy the weak global gradient domination property with $\beta=\frac{p-1}{p}$. We refer to Fatkhullin et al. [11, Appendix A] for a longer list of globally gradient dominated functions including convex and non-convex functions. Notably, in reinforcement learning it is known that the tabular softmax parametrization leads to a parametrized value function that satisfies the so-called ”non-uniform” PL-inequality [36, 37, 22]. In Section 7 we will show how this non-uniform gradient domination implies local gradient domination in $f^{*}$. This renders our local analysis of stochastic gradient methods specifically applicable in RL. As mentioned earlier, since every analytic function already satisfies local gradient domination, we expect that the local analysis can encompass further parametrizations, such as neural networks.
2.1 Assumptions on the Stochastic First Order Oracle
@@
It is worth noting that the Equation ABC assumption is a generalization of the bounded variance assumption that appears for $A=B=0$. It was introduced by Khaled and Richtárik [21] as expected smoothness condition and shown to be the weakest assumption among many others.
2.2 Stochastic Gradient Methods
The following two classical optimization algorithms will be analyzed in this article. Both algorithms are described as discrete time stochastic process $(X_{n})_{n\in\mathbb{N}}$ driven by noisy gradient evaluations in Equation 5. In each iteration, we assume that the stochastic first order oracle is accessed through the evaluation of $\zeta_{n+1}$ which is a copy of $\zeta$ independent from the current state $X_{n}$.